OpenStudy (anonymous):

The sum of a number and twice a second number is 146. The difference of the first number and second number is 5. Find the two numbers.

OpenStudy (ranga):

Let the first number be A and the second number be B. "The sum of a number and twice a second number is 146": A + 2*B = 146 ---- (1) "The difference of the first number and second number is 5": A - B = 5 ------ (2) 2 equations and 2 unknowns A & B. Solve equations (1) and (2).

OpenStudy (anonymous):

can you solve it @ranga

OpenStudy (ranga):

A + 2*B = 146 ---- (1) A - B = 5 ---- (2) Subtract (2) from (1) to get rid of A:

OpenStudy (ranga):

A + 2*B = 146 ---- (1) A - B = 5 ---- (2) ----------------------- Subtract: 0 + 3B = 141 3B = 141 B = 141/3 = 47 Put B = 47 in (2): A - 47 = 5 A = 5 + 47 = 52 The numbers are: 52 and 47.